Factor the trinomial 4x2 + 24x + 11 on your paper. Which of the following would be part of the factored answer?

Found 2 solutions by stanbon, funmath:
Answer by stanbon(29603)

Factor the trinomial 4x^2 + 24x + 11
Think of two numbers whose product is 4*11 = 44
and whose sum is 24.
The numbers are 22 and 2.
So, Replace 24x by 22x+2x to get:
=4x^2+22x+2x+11
Factor the 1st two and the last two terms separately:
=2x(2x+11)+1(2x+11)
Factor again to get:
=(2x+11)(2x+1)
======================
Cheers,
Stan H.

Answer by funmath(2873)

Factor the trinomial 4x2 + 24x + 11 on your paper. Which of the following would be part of the factored answer?
If you multiply the the coefficient of x^2 and the constant you get: 4*11=44
We need to replace the coefficient of x with two coefficients that multiply to give you 44, and add to give you the coefficient of x: 24
22*2=44 and 22+2=24
4x^2+22x+2x+11 Factor by grouping:
(4x^2+22x)+(2x+11)
2x(2x+11)+1(2x+11)
(2x+1)(2x+11)
Happy Calculating!!!